# BalanceKV: KV Cache Compression through Discrepancy Theory

### v1

@@ Abstract
Large language models (LLMs) have achieved impressive success, but their high memory requirements present challenges for long-context token generation. The memory complexity of long-context LLMs is primarily due to the need to store Key-Value (KV) embeddings in their KV cache. We present BalanceKV, a KV cache compression method based on geometric sampling process stemming from Banaszczyk’s vector balancing theory, which introduces dependencies informed by the geometry of keys and value tokens, and improves precision. BalanceKV offers both theoretically proven and empirically validated performance improvements over existing methods.
@@ 1 Introduction
Transformer-based models are the foundation of ongoing artificial intelligence revolution. Their applications span a wide range of domains, from leading-edge language models (LLM) [AAA+23, Ant24] to text-to-image [RDN+22, Fir23, Mid22], text-to-video synthesis [Ope24b], coding assistant [Cop23] and even multimodal domain across text, audio, image, and video [Ope24a]. At the core of these models is the Transformer architecture, powered by the self-attention mechanism [VSP+17], which enables effective capture of pairwise correlations across tokens in an input sequence. As these models scale in size and context length [KMH+20], they face significant computational challenges, particularly in terms of memory usage.
@@
To unlock the full potential of LLMs, it is essential to achieve both efficiency and accuracy in generating long sequences. Most large language models, along with multimodal and video models, adopt an autoregressive, decoder-only architecture. This architecture generates tokens sequentially, applying attention dynamically to each newly generated token. To avoid redundant attention score computations during the generation phase, these models maintain a KV cache, which stores the key and value embeddings of previously generated tokens in each attention layer. However, deploying autoregressive Transformers is computationally demanding. A major challenge lies in the substantial memory requirements, as the KV cache size scales with both the model size (i.e., the number of layers and attention heads) and, critically, the context size. Additionally, each model session typically requires its own dedicated KV cache, further exacerbating memory usage.
This growing demand has become a significant bottleneck, affecting both memory consumption and computational speed, particularly for models handling long context lengths. Therefore, reducing the KV cache size without compromising accuracy is essential to overcoming these limitations.
Several approaches have been proposed to address this challenge. Architectural solutions, such as multi-query attention [Sha19], multi-group attention [ALTdJ+23], and sparse mixture of experts [DDZ+24], require modifying the model during the training stage and retraining it from scratch. As a result, these methods are not directly applicable as off-the-shelf solutions for reducing the KV cache without incurring substantial pretraining costs. Another line of research tackles the issue from a systems perspective, such as RAM offloading [SZY+23, SCB+24] or integrating virtual memory and paging strategies into the attention mechanism [KLZ+23].
In this work, we propose BalanceKV, a novel approach based on discrepancy theory designed to significantly reduce the runtime and memory complexity for KV cache. Our method can shift conventional linear scaling to a sublinear scale with respect to context length. The core of our approach is a vector balancing algorithm from discrepancy theory that exploits the geometry of key and value tokens to deduce a small subset of them that well approximates the operations happening inside a self-attention layer.
1.1 Related Works
The direct approach to KV caching stores all key and value representations for every input token, leading to significant memory growth in long sequences, but several algorithmic ideas for improvement have been developed in the literature.
Embedding quantization.
A simple yet effective approach to reducing KV cache size is quantizing the floating-point numbers (FPN) in the KV cache to fixed-point numbers with fewer bits. Several quantization methods have been proposed specifically for the KV cache [YYD+24, YKB+24, DCQW24, KZK+24, LYJ+24, HKM+24, ZYXS24, ZDH24]. However, such approaches still require memory space that linearly increases in sequence length.
Token level pruning.
A prominent line of work which is related to our paper focuses on token-level KV cache pruning where redundant or less important tokens get evicted from the cache [BPC20, ZSZ+24, LDL+24, XTC+23, ZHMK24, LHY+24]. Many of the works in this line have used accumulated attention scores to select important tokens in the cache [ZSZ+24, LHY+24, XTC+23]. Recent works extend those methods to an adaptive way of budget allocation across layer [CZG+24] and head [FCA+24].
A very recent work, Lexico [KPCP24], applies techniques from sparse representation learning to compress the KV cache by learning a universal dictionary such that all key and value embeddings are represented as extremely sparse vectors within the learned dictionary. Unfortunately, this approach requires solving a computationally expensive matching pursuit algorithm for each key and value embedding, making Lexico relatively slow.
Algorithmic discrepancy theory.
@@ 1.1 Related Works
Banaszczyk’s seminal works [Ban98, Ban12] establishing theoretical guarantees for vector set discrepancy have sparked research in the vector balancing problem [DNTTJ18]. This led to algorithmic developments in both offline [Ban10] and online [BJSS19, ALS21, KRR23] settings. The vector balancing problem has particular relevance to streaming and sublinear algorithms, as minimizing a dataset’s discrepancy yields small subsets that effectively preserve the original dataset’s properties. Recent works [PT20, CKW24] extend these discrepancy theory ideas to develop sublinear memory algorithms for kernel density estimation.
@@ 1.2 Overview of Our Contributions
For a sequence of tokens and their corresponding query, key and value embeddings in an input prompt, a natural approach to compressing the exact KV cache would be to first select key and value embeddings for a few important tokens in the sequence such as the first and latest few tokens, as motivated by [XTC+23], since important contexts are likely contained in them, and then select a small representative subset of key and value embeddings for the much longer middle portion of the sequence. This representative subset can be chosen by independently sampling and selecting key and value embeddings from middle portion, however this would ignore the semantic relationship of the tokens in the middle portion encoded in the geometric structure of the corresponding key and value embeddings. Several powerful practical heuristic approaches, discussed in Section 1.1, have been proposed recently, PyramidKV [CZG+24] and SnapKV [LHY+24] being the state of the art. These heuristics select a subset of key and value embeddings from the entire sequence using attention score computation, thereby taking relative semantic importance of input tokens into account. However, these techniques lack theoretical guarantees. Our main contribution is a principled geometric sampling method for selecting a representative subset of tokens, which matches or outperforms the above mentioned heuristics. Our contributions are as follows:
@@
    In Section 3 we propose BalanceKV, an algorithm for compressing the KV cache recursively using a geometric correlated sampling process based on discrepancy theory. We show that BalanceKV provably approximates attention in the streaming setting under the bounded $\ell_{2}$ norm assumption (Theorem 3.1). Section 2 contains the formal problem formulation of streaming attention as well as a technical overview of the main results and techniques.
@@
    In Section 4 we empirically evaluate our algorithm in two settings. In the first setting (see Section 4.1) we show our approach leads to a lower relative error for single layer attention approximation for open-source LLMs including $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$ [DJP+24] and $\mathtt{Ministral}$-$\mathtt{8B}$-$\mathtt{Instruct}$-$\mathtt{2410}$ [Mis24] as compared to uniformly sampling keys and values in the cache. In the second setting we show that our provable method performs better compared to previous existing heuristic cache compression methods based on selecting a subset of key and value embeddings in the cache on the LongBench [BLZ+23] benchmark for long context understanding capabilities of large language models under various end to end tasks.
@@ 2 Technical Overview
2.1 KV Cache Compression
@@
Formally, suppose we have a stream of query, key and value embeddings $(q_{1},k_{1},v_{1}),\ldots,(q_{n},k_{n},v_{n})$, that is the $j$-th token is represented as a triplet of ($q_{j},k_{j},v_{j}$) where $q_{j},k_{j},v_{j}\in\mathbb{R}^{d}$ for all $j\in[n]$. Let $K_{j},V_{j}\in\mathbb{R}^{j\times d}$ be matrices defined by stacking those keys and values in their respective rows. To compute the following at every step $j$ to generate $j+1$ token, is called the streaming attention problem:
@@
Keeping all of the key-value pairs in the cache is prohibitively expensive, especially for long sequences. Instead, we opt for approximate computation by sampling a few key-value pairs. Specifically, our goal is to construct an algorithm that at every time step $j$ computes an estimator $z_{j}$ for $\text{Attn}(q_{j},K_{j},V_{j})$ in sublinear in $n$ time and memory. In particular for given precision $\varepsilon$, $z_{j}$ should satisfy the following error constraint:
@@
A sublinear in $n$ time and memory algorithm to compute $z_{j}$ will require knowledge of significantly less key-value pairs than $K_{j},V_{j}$, thus reducing the size of the KV cache needed to store them. In the next section we discuss how we will construct such an estimator $z_{j}$ at a high level.
@@ 2.2 SoftmaxBalance: Attention Approximation via Discrepancy Theory
where for a matrix $A$ we write $\exp(A)$ to denote entry-wise application of the exponential function to $A$ and
$Z_{j}=\sum_{i\in[j]}\exp\left(\frac{\langle k_{i},q_{j}\rangle}{\sqrt{d}}\right).$
Our approach to approximate Attn$(q_{j},K_{j},V_{j})$ consists of two subroutines which approximate:
  1. 1.
    Softmax normalization $Z_{j}=\sum_{i\in[j]}\exp\left(\frac{\langle k_{i},q_{j}\rangle}{\sqrt{d}}\right)$,
  2. 2.
    Matrix-vector product between $V_{j}$ and $\exp\left(\frac{K_{j}\cdot q_{j}}{\sqrt{d}}\right)$.
To understand our main idea, suppose we are at the end of the stream (i.e., $j=n$) and the KV cache contains all key-value pairs $(k_{1},v_{1}),\ldots,(k_{n},v_{n})$. Then for an arbitrary query $q_{n}$ we aim to approximate the matrix-vector product
$\exp\left(\frac{K_{n}\cdot q_{n}}{\sqrt{d}}\right)^{T}\cdot V_{n}=\sum_{i\in[n]}\exp\left(\frac{\langle k_{i},q_{n}\rangle}{\sqrt{d}}\right)v_{i}$
by choosing a subset of the rows of $K_{n}$ and $V_{n}$ of size at most $n/2$ which corresponds to a compression rate of $0.5$. Suppose we can design an algorithm which splits the set $C$ of all keys and values into two groups $C^{\prime}$ and $C\backslash C^{\prime}$ so that the matrix-vector product function for any query vector $q_{n}$ is roughly equal over $C^{\prime}$ and $C\backslash C^{\prime}$ that is informally,
@@
Therefore, it would suffice to keep the smaller subset of $C^{\prime}$ and $C\backslash C^{\prime}$ as the desired subset of key value embeddings and discard the rest. If we wanted to compress the KV-cache to a smaller size by a factor $2^{T}$ for some $T$, we would recursively compress the selected subset using the same procedure $T-1$ more times.
@@
A similar goal is captured by the vector balancing problem studied extensively in discrepancy theory - given a set of vectors $C=\{k_{1},\dots,k_{n}\}\subset\mathbb{R}^{d}$ with $\|k_{j}\|_{2}\leq 1$ for all $j$, partition them into two groups $C^{\prime},C\setminus C^{\prime}$ such that for any $q\in\mathbb{R}^{d}$ it holds $\sum_{k\in C^{\prime}}\langle k,q\rangle\approx\sum_{k\in C\setminus C^{\prime}}\langle k,q\rangle$ with high probability. The Self-Balancing Walk algorithm [ALS21] is a breakthrough result which gives an algorithm for the above vector balancing problem. However we need to develop an algorithm for the vector balancing problem with respect to function $\exp\left(\langle k,\cdot\rangle/\sqrt{d}\right)v$ instead of the inner product function $\langle k,\cdot\rangle$.
@@
Our first contribution is to develop an algorithm for our task, building upon the result from [ALS21], which essentially randomly partitions the set of keys and values $C$ into $C^{\prime}$ and $C\setminus C^{\prime}$ such that the following holds with high probability under the assumptions that the norms of the query and key embeddings are bounded,
@@
$\left\|\sum_{\{k,v\}\in C^{\prime}}\exp\left(\frac{\langle k,q_{n}\rangle}{\sqrt{d}}\right)v-\sum_{\{k,v\}\notin C^{\prime}}\exp\left(\frac{\langle k,q_{n}\rangle}{\sqrt{d}}\right)v\right\|_{2}\leq O\left(\log(nd)\right)\cdot\max_{j\in[n]}\|v_{j}\|_{2}.$
@@
We refer to this algorithm as SoftmaxBalance, its formal guarantees are presented in Theorem 3.2 and its pseudocode is presented in Algorithm 2. Theorem 3.2 shows that SoftmaxBalance succeeds to divide $C$ into subsets $C^{\prime}$ and $C\backslash C^{\prime}$ which are balanced with respect to function $\exp\left(\langle k,\cdot\rangle/\sqrt{d}\right)v$ up to an error which only has logarithmic dependence on the size of $C$. In addition, SoftmaxBalance can accept as input value vectors of arbitrary dimension $s$. Therefore, if instead of the value vectors $V=\{v_{1},\ldots,v_{n}\}\subset\mathbb{R}^{d}$ we input the set of scalars $V=\{1,\ldots,1\}$, we will get an algorithm for the vector balancing problem with respect to function $\exp\left(\langle k,\cdot\rangle/\sqrt{d}\right)$. This implies that we can use SoftmaxBalance to compress the KV cache to even approximate the softmax normalization $\sum_{i\in[n]}\exp\left(\langle k_{i},q_{n}\rangle/{\sqrt{d}}\right)$.
We now discuss how to use SoftmaxBalance for streaming attention approximation, that is to use its guarantees to compute an estimator $z_{j}$ satisfying Equation 2.
2.3 BalanceKV: Implementing SoftmaxBalance in Streaming for Cache Compression

### v2

# Streaming Attention Approximation via Discrepancy Theory

@@ Abstract
Large language models (LLMs) have achieved impressive success, but their high memory requirements present challenges for long-context token generation. In this paper we study the streaming complexity of attention approximation, a key computational primitive underlying token generation.
Our main contribution is BalanceKV, a streaming algorithm for $\epsilon$-approximating attention computations based on geometric process for selecting a balanced collection of Key and Value tokens as per Banaszczyk’s vector balancing theory. We complement our algorithm with space lower bounds for streaming attention computation. Besides strong theoretical guarantees, BalanceKV exhibits empirically validated performance improvements over existing methods, both for attention approximation and end-to-end performance on various long context benchmarks.
@@ 1 Introduction
Transformer-based models are the foundation of ongoing artificial intelligence revolution. Their applications span a wide range of domains, from leading-edge language models (LLM) [AAA+23, Tea24a] to text-to-image [RDN+22, Tea23a, Tea22], text-to-video synthesis [Tea24c], coding assistance [Tea23b] and even in multimodal domains across text, audio, image, and video [Ope24]. At the core of these models is the Transformer architecture, powered by the self-attention mechanism [VSP+17], which enables effective capture of pairwise correlations across tokens in an input sequence. As these models scale in size and context length [KMH+20], they face significant computational challenges, particularly in terms of memory usage. Efficiency and accuracy are essential to unlock the full potential of LLMs in generating long sequences.
Space bottlenecks in transformer models.
@@
Most large language models, along with multimodal and video models, adopt an autoregressive, decoder-only architecture. This architecture generates tokens sequentially, applying attention dynamically to each newly generated token. To avoid redundant attention score computations during the generation phase, these models explicitly store the key and value embeddings of previously generated tokens in a cache in each attention layer. Thus, a major challenge is the fact that the memory complexity of storing previously generated key value embeddings scales with both the model size (i.e., the number of layers and attention heads) and, critically, the context size. Additionally, each model session typically requires its own dedicated cache for storing key value embeddings, further exacerbating memory usage. This growing demand has become a significant bottleneck, affecting both memory consumption and computational speed, particularly for models handling long context lengths.
Streaming attention computation.
The main reason for the need of storing the past key and value embeddings is for the attention computation happening inside each self attention layer during token generation after processing a context – to generate the next token, each self attention layer computes the attention between the query embedding of the current token and the key and value embeddings of all the tokens that were previously generated or part of the context. In this paper we study the streaming attention approximation problem – the problem of approximately computing attention using a small amount of space, i.e. without storing all previously seen key and value embeddings. Our main contribution is BalanceKV, a novel provably correct algorithm for streaming attention approximation based on discrepancy theory. The core of our approach is a vector balancing algorithm from discrepancy theory that exploits the geometry of key and value tokens to deduce a small subset of them that well approximates the operations happening inside a self-attention layer. We complement our algorithm with a lower bound on the streaming complexity of approximating attention.
An algorithm for streaming attention approximation can directly be used for compressing the key value cache which stores the past key value embeddings in each layer in an LLM, thus improving the efficiency of LLM token generation. We empirically evaluate BalanceKV both on the problem of approximating attention and on end-to-end generation tasks, showing performance gains.
1.1 Related Work
@@
For discrepancy theory, Banaszczyk’s seminal works [Ban98, Ban12] establishing theoretical guarantees for vector set discrepancy have sparked research in the vector balancing problem [DNTTJ18]. This led to algorithmic developments in both offline [Ban10] and online [BJSS19, ALS21, KRR23] settings. The vector balancing problem has particular relevance to streaming and sublinear algorithms, as minimizing a dataset’s discrepancy yields small subsets that effectively preserve the original dataset’s properties. Recently  [PT20, CKW24] extend these discrepancy theory ideas for kernel density estimation using sublinear memory.
A simple yet effective approach is quantizing previously generated key value embeddings with fewer bits  [YYD+24, YKB+24, DCQW24, KZK+24, LYJ+24, HKM+24, ZYXS24, ZDH24].
Another line of work focuses on token-level pruning, where redundant or less important tokens get evicted from the set of all previously generated key value embeddings [BPC20, ZSZ+24, LDL+24, XTC+23, ZHMK24, LHY+24]. Many of the works in this line have used accumulated attention scores to select important previously generated tokens [ZSZ+24, LHY+24, XTC+23]. Recent works extend those methods to an adaptive way of budget allocation across layer [CZG+24] and head [FCA+24].
@@ 1.2 Overview of Our Contributions
In this work we take the token subset selection approach to reduce the memory complexity of LLM token generation: store and maintain only a subset of previously generated key and value embeddings corresponding to a few “important” tokens in the sequence. Of course, the central question is how to define “importance” of tokens. Our approach here is to apply discrepancy theory, which, at a high level, considers a token important if it is crucial to preserving the projection of the total collection of tokens onto some direction in the token space. This leads to the idea of selecting a subset of tokens that is “balanced” simultaneously in every direction. Inspired by the recent breakthrough result of [ALS21] on online discrepancy minimization, we design a method for balancing key-value pairs online using small space, namely our BalanceKV algorithm. Interestingly, this algorithm is online, i.e. the importance of a token is determined only by preceding tokens – in sharp contrast with state of the art heuristics for token selection such as PyramidKV [CZG+24] and SnapKV [LHY+24], whose performance, as we show, our algorithm matches or improves upon. Our contributions are:
@@
    In Section 3 we propose BalanceKV, an algorithm for recursively compressing the set of previously generated tokens using a geometric correlated sampling process based on discrepancy theory. We show that BalanceKV gives provable guarantees for streaming attention approximation under the bounded $\ell_{2}$ norm assumption (Theorem 3.1). Using tools from communication complexity, we also show a lower bound on the memory complexity of any algorithm for streaming attention approximation in Section 3.
    Section 2 contains the formal problem formulation of streaming attention approximation, its applicability to key value cache compression, as well as a technical overview of the main results and techniques of Section 3.
@@
    In Section 4 we empirically evaluate our algorithm in various settings. In Section 4.1 we show our approach leads to a lower relative error for single layer attention approximation for open-source LLMs including $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$ [DJP+24] and $\mathtt{Ministral}$-$\mathtt{8B}$-$\mathtt{Instruct}$-$\mathtt{2410}$ [Mis24] as compared to uniformly sampling keys and values in the cache. Section 4.1 we also perform ablation studies to show how various parameters in our algorithm affect the relative error for single layer attention approximation. In Sections 4.2 and 4.3 we perform end to end experiments on various benchmarks such as LongBench [BLZ+23] using models of various sizes such as $\mathtt{Llama}$-$\mathtt{3.1}$-$\mathtt{8B}$-$\mathtt{Instruct}$, $\mathtt{Qwen}$-$\mathtt{2.5}$-$\mathtt{14B}$-$\mathtt{Instruct}$ and $\mathtt{Qwen}$-$\mathtt{2.5}$-$\mathtt{32B}$-$\mathtt{Instruct}$ [YYH+24, Tea24d], and Needle in a Haystack [Kam23]. We show that our provable method for attention approximation when applied to key value cache compression performs better compared to previous existing token subset selection heuristics on end to end tasks. Finally in Section 4.4 we present system efficiency metrics regarding our implementation.
@@ 2 Technical Overview
2.1 Streaming Attention Approximation: Formulation and Motivation
@@
Formally, suppose that we have a stream of query, key and value embeddings $(q_{1},k_{1},v_{1}),\ldots,(q_{n},k_{n},v_{n})$, that is the $j$-th token is represented as a triplet of ($q_{j},k_{j},v_{j}$) where $q_{j},k_{j},v_{j}\in\mathbb{R}^{d}$ for all $j\in[n]$. Let $K_{j},V_{j}\in\mathbb{R}^{j\times d}$ be matrices defined by stacking those keys and values in their respective rows.
To compute the following at every step $j$ to generate $j+1$ token, is called the streaming attention problem:
@@
Keeping all of the key-value pairs in the cache is prohibitively expensive, especially for long sequences. Instead, we opt for approximate computation by sampling a few key-value pairs. Specifically, our goal is to construct an algorithm that at every time step $j$ computes an estimator $z_{j}$ for $\text{Attn}(q_{j},K_{j},V_{j})$ in sublinear in $n$ time and memory. In particular for given precision $\varepsilon>0$, $z_{j}$ should satisfy the following error constraint:
@@
A sublinear in $n$ time and memory algorithm to compute $z_{j}$ will require knowledge of significantly less key-value pairs than $K_{j},V_{j}$, thus reducing the size of the key value cache needed to store them. This motivates the study of streaming attention approximation, as an algorithm for this can directly be used for key value cache compression during LLM token generation. In the next section we discuss how we will construct such an estimator $z_{j}$ at a high level.
@@ 2.2 SoftmaxBalance: Attention Approximation via Discrepancy Theory
where for a matrix $A$ we write $\exp(A)$ to denote entry-wise exponential function to $A$ and
$\ Z_{j}:=\sum_{i\in[j]}\exp(\langle k_{i},q_{j}\rangle/\sqrt{d}).$
Our approach to approximate Attn$(q_{j},K_{j},V_{j})$ consists of two subroutines which approximate:$\bf(1)$ Softmax normalization $Z_{j}=\sum_{i\in[j]}\exp(\langle k_{i},q_{j}\rangle/{\sqrt{d}})$, $\bf(2)$ Matrix-vector product between $V_{j}$ and $\exp(K_{j}\cdot q_{j}/\sqrt{d})$. To understand our main idea, suppose we are at the end of the stream (i.e., $j=n$) and we store all key-value pairs $(k_{1},v_{1}),\ldots,(k_{n},v_{n})$. Then for an arbitrary query $q_{n}$ we aim to approximate the matrix-vector product $\exp(K_{n}\cdot q_{n}/\sqrt{d})^{T}\cdot V_{n}=\sum_{i\in[n]}\exp(\langle k_{i},q_{n}\rangle/\sqrt{d})v_{i}$ by choosing a subset of the rows of $K_{n}$ and $V_{n}$ of size at most $n/2$ which corresponds to a compression rate of $0.5$. Suppose we can design an algorithm which splits the set $C$ of all keys and values into two groups $C^{\prime}$ and $C\backslash C^{\prime}$ so that the matrix-vector product function for any query vector $q_{n}$ is roughly equal over $C^{\prime}$ and $C\backslash C^{\prime}$ that is informally,
@@
Therefore, it would suffice to keep the smaller subset of $C^{\prime}$ and $C\backslash C^{\prime}$ as the desired subset of key value embeddings and discard the rest. If we wanted to compress the key value cache to a smaller size by a factor $2^{T}$ for some $T$, we would recursively compress the selected subset using the same procedure $T-1$ more times.
@@
A similar goal is captured by the vector balancing problem studied extensively in discrepancy theory; given a set of vectors $C=\{k_{1},\dots,k_{n}\}\subset\mathbb{R}^{d}$ with $\|k_{j}\|_{2}\leq 1$ for all $j$, partition them into two groups $C^{\prime},C\setminus C^{\prime}$ such that for any $q\in\mathbb{R}^{d}$ it holds $\sum_{k\in C^{\prime}}\langle k,q\rangle\approx\sum_{k\in C\setminus C^{\prime}}\langle k,q\rangle$ with high probability. The Self-Balancing Walk algorithm [ALS21] is a breakthrough result for the above vector balancing problem. However we need to develop an algorithm for the vector balancing problem with respect to function $\exp(\langle k,\cdot\rangle/\sqrt{d})v$ instead of the inner product function $\langle k,\cdot\rangle$.
@@
Our first contribution is to develop an algorithm for our task, building upon the result from the self-balancing walk [ALS21], which essentially randomly partitions the set of keys and values $C$ into $C^{\prime}$ and $C\setminus C^{\prime}$ such that the following holds with high probability under the assumptions that the norms of the query and key embeddings are bounded,
@@
$\left\|\sum_{\{k,v\}\in C^{\prime}}\exp\left(\frac{\langle k,q_{n}\rangle}{\sqrt{d}}\right)v-\sum_{\{k,v\}\notin C^{\prime}}\exp\left(\frac{\langle k,q_{n}\rangle}{\sqrt{d}}\right)v\right\|_{2}\leq O\left(\log(nd)\right)\cdot\max_{j\in[n]}\|v_{i}\|_{2}.$
@@
We refer to this algorithm as SoftmaxBalance, its formal guarantee is presented in Theorem 3.2 and its pseudocode is presented in Algorithm 2. Theorem 3.2 shows that SoftmaxBalance succeeds to divide $C$ into subsets $C^{\prime}$ and $C\backslash C^{\prime}$ which are balanced with respect to function $\exp(\langle k,\cdot\rangle/\sqrt{d})v$ up to an error which only has logarithmic dependence on the size of $C$. In addition, SoftmaxBalance can accept as input value vectors of arbitrary dimension $s$. Therefore, if instead of the value vectors $v_{1},\ldots,v_{n}\in\mathbb{R}^{d}$ we input the set of scalars $v_{1}=\dots=v_{n}=1$, we will get an algorithm for the vector balancing problem with respect to function $\exp(\langle k,\cdot\rangle/\sqrt{d})$. This implies that we can use SoftmaxBalance to compress the key value cache to even approximate the softmax normalization $\sum_{i\in[n]}\exp(\langle k_{i},q_{n}\rangle/{\sqrt{d}})$. We now discuss how to use SoftmaxBalance for streaming attention approximation, i.e. to use it to compute an estimator $z_{j}$ satisfying Equation 2.
2.3 BalanceKV: Implementing SoftmaxBalance in Streaming